A rhombus is formed by two radii and two chords of a circle of radius 10 m. What is the area of  the rhombus?

Grade:12th pass

1 Answers

Harshit Singh
3 years ago
Dear Student

As diagonals of rhombus perpendicularly bisect each other.
So,OB=10 m
OP=5 cm
LetPC=AP=x
Now,
OC2=OP2+PC2
102=52+x2
x2=75
x=5√3​ m
AC=10√3​ m
Area of rhombus = 10×10√​3/2=50√3 sq. m.

Thanks